# SysCaps: Language Interfaces for Simulation Surrogates of Complex Systems

### v1

@@ Appendix A Additional Experiment Details
$:=\frac{1}{\sum y_{t}^{b}}\sqrt{((\sum_{b=1,t=1}^{B,T}y^{b}_{t})-(\sum_{b=1,t=1}^{B,T}\hat{y}^{b}))^{2}}.$ (3)
We use the AdamW [35] optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.98, $\epsilon$ = 1e-9, and weight decay of 0.01 for all experiments. The early stopping patience is 50 for all experiments. All models are trained with a single NVIDIA A100-40GB GPU. The longest training runs take 1-2 days and the shortest 2-3 hours.
A.1 Hyperparameter sweeps
See Table 5 and Table 6.
A.2 Buildings
@@
Model Hyperparameter Grid search space Best values
LightGBM Learning rate From 0.01 to 0.1 0.066
Number of leaves From 40 to 150 149
Subsample From 0.05 to 1.0 0.178
Feature fraction From 0.05 to 1.0 0.860
Min number of data in one leaf From 1 to 100 12
ResNet (one-hot) Hidden layers size 256, 1024, 2048 1024
Number of layers 2, 8 2
Batch size 128, 256, 512 512
Learning rate 0.0001, 0.0003, 0.001 0.0003
ResNet (SysCaps) Hidden layers size 256, 1024, 2048 256
Number of layers 2, 8 8
Batch size 128, 256, 512 256
Learning rate 0.0001, 0.0003, 0.001 0.0003
Bidirectional LSTM (one-hot) [Hidden layer size, Batch size] [128, 64], [512, 32], [1024, 32] [128, 64]
Number of layers 1, 3, 4, 6, 8 4
MLP dimension 256 256
Learning rate 0.00001, 0.0003, 0.001 0.001
Bidirectional LSTM (SysCaps) [Hidden layer size, Batch size] [128, 64], [512, 32], [1024, 32] [512, 32]
Number of layers 1, 3, 4, 6, 8 1
MLP dimension 256 256
Learning rate 0.00001, 0.0003, 0.001 0.0003
Bidirectional S4 (one-hot) [Hidden layer size, Num. layers] [64,8] , [128,4] [128,4]
MLP dimension 256 256
Batch size 32, 64 64
Learning rate 1e-5, 3e-4, 1e-3 3e-4
Bidirectional S4 (SysCaps) [Hidden layer size, Num. layers] [64,8] , [128,4] [128, 4]
MLP dimension 256 256
Batch size 32, 64 32
Learning rate 1e-5, 3e-4, 1e-3 3e-4
There are 13 attributes after RFE, which are one-hot encoded into a 336-dimensional feature vector, whereas the text embeddings are 768-dimensional. Along with the 7 weather variables, we concatenate cyclically encoded calendar features [15]. This creates a 103-dimensional input for the bidirectional sequence encoders.
LightGBM: As LightGBM does not support batch training out of the box, the entire training data needs to be loaded into the memory to train a LightGBM model. With the train dataset containing 340k buildings, each with 8759 hours and 347 features, we randomly extract 438 hours per building (which is about 5% of total hours) to limit memory usage. This results in $340,000\times 438$ hours in total for the train dataset, which consumes about 380 GB of memory when being loaded into a NumPy object. For the validation and test splits, we retain the full number of hours per building. The LightGBM model is tuned with Optuna [2] across 30 trials and achieves the best validation NRMSE of 0.667.
A.2.1 Wind farm
@@
LightGBM: The training, validation, and test split for the wind dataset gives us datasets of size 148,650, 49,600, and 49,250 respectively with 190 features after one-hot encoding. The training dataset is loaded into memory to train the LightGBM model. We use Optuna [2] to tune the hyperparameters and the best validation NMRSE achieved is 0.189.
@@ Appendix B Additional SysCaps examples
Figure 7: Building SysCap.
Figure 8: Wind farm SysCap.
Figure 9: SysCap building with an incorrectly described attribute (weekday closing time - highlighted in blue) in the natural language caption.
Figure 10: Wind farm SysCap with a logical error in the natural language caption where it says "total installed capacity" when it is just the capacity of a single turbine (highlighted in blue).

### v2

@@ A.1 Metrics
$:=\frac{1}{\sum y_{t}^{b}}\sqrt{\bigg{(}(\sum_{b=1,t=1}^{B,T}y^{b}_{t})-(\sum_{b=1,t=1}^{B,T}\hat{y}^{b})\bigg{)}^{2}}.$ (3)
Notice that the square and square root cancel, making the stock-annual NRMSE equivalent to the normalized mean bias error.
We use the AdamW Loshchilov & Hutter (2017) optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.98, $\epsilon$ = 1e-9, and weight decay of 0.01 for all experiments. The early stopping patience is 50 for all experiments. All models are trained with a single NVIDIA A100-40GB GPU. The longest training runs take 1-2 days and the shortest 2-3 hours.
A.2 Hyperparameters
@@
Model Hyperparameter Grid search space Best values LightGBM Learning rate From 0.01 to 0.1 0.066 Number of leaves From 40 to 150 149 Subsample From 0.05 to 1.0 0.178 Feature fraction From 0.05 to 1.0 0.860 Min number of data in one leaf From 1 to 100 12 ResNet (one-hot) Hidden layers size 256, 1024, 2048 1024 Number of layers 2, 8 2 Batch size 128, 256, 512 512 Learning rate 0.0001, 0.0003, 0.001 0.0003 ResNet (SysCaps) Hidden layers size 256, 1024, 2048 256 Number of layers 2, 8 8 Batch size 128, 256, 512 256 Learning rate 0.0001, 0.0003, 0.001 0.0003 Bidirectional LSTM (one-hot) [Hidden layer size, Batch size] [128, 64], [512, 32], [1024, 32] [128, 64] Number of layers 1, 3, 4, 6, 8 4 MLP dimension 256 256 Learning rate 0.00001, 0.0003, 0.001 0.001 Bidirectional LSTM (SysCaps) [Hidden layer size, Batch size] [128, 64], [512, 32], [1024, 32] [512, 32] Number of layers 1, 3, 4, 6, 8 1 MLP dimension 256 256 Learning rate 0.00001, 0.0003, 0.001 0.0003 Bidirectional S4 (one-hot) [Hidden layer size, Num. layers] [64,8] , [128,4] [128,4] MLP dimension 256 256 Batch size 32, 64 64 Learning rate 1e-5, 3e-4, 1e-3 3e-4 Bidirectional S4 (SysCaps) [Hidden layer size, Num. layers] [64,8] , [128,4] [128, 4] MLP dimension 256 256 Batch size 32, 64 32 Learning rate 1e-5, 3e-4, 1e-3 3e-4
@@
See Table 6 and Table 7.
A.2.1 Buildings
Figure 5: A visualization of the sequential surrogate model baseline with one-hot encoded system attributes for the buildings experiment.
There are 13 attributes after RFE, which are one-hot encoded into a 336-dimensional feature vector that gets embedded into 128 dimensions, whereas the text embeddings are 768-dimensional. We concatenate cyclically encoded calendar features to 7 weather variables, creating a 103-dimensional vector (Emami et al., 2023b). For the one-hot models (Figure 5), this creates a 128 + 103 = 231 dimensional input for the bidirectional sequence encoder, and for the text models, it is a 768 + 103 = 871 dimensional input.
LightGBM: As LightGBM does not support batch training out of the box, the entire training data needs to be loaded into the memory to train a LightGBM model. With the train dataset containing 340k buildings, each with 8759 hours and 347 features, we randomly extract 438 hours per building (which is about 5% of total hours) to limit memory usage. This results in $340,000\times 438$ hours in total for the train dataset, which consumes about 380 GB of memory when being loaded into a NumPy object. For the validation and test splits, we retain the full number of hours per building. The LightGBM model is tuned with Optuna (Akiba et al., 2019) across 30 trials and achieves the best validation NRMSE of 0.667.
A.2.2 Wind farm
LightGBM: The training, validation, and test split for the wind dataset gives us datasets of size 148,650, 49,600, and 49,250 respectively with 190 features after one-hot encoding. The training dataset is loaded into memory to train the LightGBM model. We use Optuna (Akiba et al., 2019) to tune the hyperparameters and the best validation NMRSE achieved is 0.189.
@@ Appendix B Additional SysCaps examples
In this section we visualize examples of prompts and the corresponding SysCaps outputs for the buildings simulator (Figures 6, 7, and 8) and the wind farm simulator (Figures 9, 10, and 11).
Figure 6: Natural language building SysCap. The instruction and list of key-value attributes (first two paragraphs) are provided to the LLM as the prompt. We observe that the LLM automatically converts numbers to more human-interpretable descriptions (e.g., the number of stories is changed from $1.0$ in the prompt to “single-story”). The LLM also succinctly (and correctly) states that “there is no difference in the temperature set points when the space is unoccupied” in the output by summarizing the “unoccupied heating/cooling temperature set point difference from occupied state: 0.0” attributes.
Figure 7: Natural language building SysCap. The LLM makes a subtle logical error here—the unoccupied heating temperature set point difference from occupied state suggests the set point should be lower by 10 degrees, not higher (highlighted in blue). Logical errors such as this may have contributed to the slightly worse accuracy of natural language SysCaps compared to key-value SysCaps. Collaborating with domain experts able to validate the LLM’s outputs is important for catching these errors. We expect that using more powerful LLMs will help reduce logical errors.
Figure 8: Natural language building SysCap. The LLM confuses the weekday closing time with the weekend open time (highlighted in blue).
Figure 9: Natural language wind farm SysCap.
Figure 10: Natural language wind farm SysCap.
Figure 11: Natural language wind farm SysCap with a logical error where it says "total installed capacity", but it just states the capacity of a single turbine (highlighted in blue). The correct total installed capacity is number of turbines times 3.4, or 431.8 MW.